# Large-scale semi-organized rolls in a sheared convective turbulence: Mean-field simulations

### v1

@@ V Conclusions
Mean-field simulations based on a developed mean-field theory [50, 51] of a non-rotating sheared turbulent convection are performed. This mean-field theory describes an effect of modification of the turbulent heat flux by the non-uniform large-scale motions caused by production of essentially anisotropic velocity fluctuations generated by tangling of the mean-velocity gradients by small-scale turbulent motions. The effect causes an excitation of large-scale convective-shear instability and the formation of large-scale convective rolls. During the nonlinear stage of the convective-shear instability, there is a transition from the two-layer vertical structure with two roles in the vertical direction before the system reaches steady-state to the one-layer vertical structure with one role after the system reaches steady-state. This effect is observed for all effective Rayleigh numbers.
@@ Acknowledgements.
I.R. would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge University, for support and hospitality during the programme ”Anti-diffusive dynamics: from sub-cellular to astrophysical scales”, where the final version of the paper was completed. The authors benefited from stimulating discussions of various aspects of turbulent convection with A. Brandenburg and P. J. Käpylä.

### v2

@@ V Conclusions
Mean-field simulations based on a developed mean-field theory [50, 51] of a non-rotating sheared turbulent convection are performed. This mean-field theory describes an effect of modification of the turbulent heat flux by the non-uniform large-scale motions caused by production of essentially anisotropic velocity fluctuations generated by tangling of the mean-velocity gradients by small-scale turbulent motions. The effect causes an excitation of large-scale convective-shear instability and the formation of large-scale convective rolls. During the nonlinear stage of the convective-shear instability, there is a transition from the two-layer vertical structure with two rolls in the vertical direction before the system reaches steady-state to the one-layer vertical structure with one roll after the system reaches steady-state. This effect is observed for all effective Rayleigh numbers.
@@ Acknowledgements.
I.R. would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme ”Anti-diffusive dynamics: from sub-cellular to astrophysical scales”, where the final version of the paper was completed. The work of N.K. was partially supported by the Russian Science Foundation (grant 21-72-20067). The authors benefited from stimulating discussions of various aspects of turbulent convection with A. Brandenburg and P. J. Käpylä.